Demand is represented by the equation, P = 80 - 0.3QD and supply by the equation P = 30 + 0.2QS.

(a) Determine the equilibrium price and quantity.
(b) What are the economic effects of a price ceiling at $41?
(c) What are the economic effects of a price ceiling at $72?
(d) What are the economic effects of a price floor at $62?
(e) What are the economic effects of a price floor at $37?
1
Expert's answer
2019-06-19T11:22:39-0400

a) 80-0.3Q=30+0.2Q; => 50=0.5Q, => Q=100, P=80-0.3*100=50

b) If P=$41, quantity demanded is (80-41)/0.3= 130, quantity supplied is (30-41)/(-0.2)=55, shortge is 130-55=75

c) If P=$72, quantity demanded is (80-72)/0.3= 27, quantity supplied is (30-72)/(-0.2)=210. But the price ceiling does not allow to increase prices, while decreasing is permitted. So eventually the equilibrium price of $50 dollars will be set

d) If P=$62, quantity demanded is (80-62)/0.3= 60, quantity supplied is (30-62)/(-0.2)=160, surplus is 160-60=100

e) If P=$37, quantity demanded is (80-37)/0.3= 143, quantity supplied is (30-37)/(-0.2)=35. But the price floor does not allow to decrease prices, while increasing is permitted. So eventually the equilibrium price of $50 dollars will be set